Entrance to a prestigious MBA program in India is determined by a national test where only the top 10% of the examinees are admitted to the program. Suppose it is known that the scores on this test are normally distributed with a mean of 420 and a standard deviation of 80. Parul Monga is trying desperately to get into this program.

What is the minimum score that she must earn to get admitted?
Answer #1

Given that,

mean = = 420

standard deviation = =80

Using standard normal table,

P(Z > z) = 10%

= 1 - P(Z < z) = 0.10

= P(Z < z ) = 1 - 0.10

= P(Z < z ) = 0.90

= P(Z < 1.28 ) = 0.90

z = 1.28 (using standard normal (Z) table )

Using z-score formula  

x = z * +

x= 1.28*80+420

x= 522.4

x=522